Thanks for the help. I've tried every equation I can think of and I haven't come up with a suitable answer.
Word Problem:
Brian's present age is one year more than twice Shelly's age. Five years ago, Brian was four times as old as Shelly. How old is Brian now?

Answer by Earlsdon(6294)

Let B = Brian's present age and S = Shelly's present age. From the problem description, we get:
1) B = 2S+1 Brian's age is one year more (+1) than twice Shelly's age (2S).
2) B-5 = 4(S-5) Five years ago (-5) Brian was 4 times Shelly's age 4(S-5)
Substitute equation 1) into equation 2) and solve for S.
(2S+1)-5 = 4(S-5) Simplify.
2S-4 = 4S-20 Subtract 2S from both sides of the equation.
-4 = 2S-20 Add 20 to both sides.
16 = 2S Divide both sides by 2.
S = 8 This is Shelly's age.
B = 2S+1
B = 2(8)+1
B = 16+1
B = 17 This is Brian's age.
Check:
B = 2S+1
17 = 2(8)+1
17 = 17
B-5 = 4(S-5)
17-5 = 4(8-5)
12 = 4(3)
12 = 12